# Probing CPT invariance with top quarks at the LHC

### v1

@@ Abstract
CPT invariance is a symmetry of the Standard Model (SM) and generic prediction of local, unitary, and Lorentz invariant quantum field theories in Minkowski spacetime [1, *Pauli:1955, *Luders:1957bpq]. Among its consequences are the equalities of particle and antiparticle masses and lifetimes.
Connecting theoretical predictions of CPT invariance violation to particle–antiparticle CPT tests has been carried out for all particle species of the SM with exception of the top quark. This work addresses this gap, establishing the first sensitivity to top-sector CPT violation from top and antitop mass reconstructions in ATLAS and CMS Collaboration measurements.
Kostelecký and Potting demonstrated that spontaneous CPT violation in string theory may result in remnant observables (likely suppressed by the energy scale of this violation) which can be possibly tested in current experiments [4, *kp95]. The subsequent development of effective field theory (EFT) descriptions [6, *ck98, *cg97, *cg98] generated intense interest in testing fundamental spacetime symmetries and have produced hundreds of constraints across a variety of systems [10]. Within this EFT framework, known as the Standard Model Extension (SME), are operators involving top-quark fields.
Lorentz- and CPT-violating effects described by the SME can be viewed as generalized tensor background fields. In view of the discovery of the Higgs boson [11, 12] described by a homogeneous and isotropic scalar background with $SU(2)\times U(1)$ quantum numbers, searching for new physics in the form of additional backgrounds is well motivated.
In this study, we suggest using measurements of the kinematically reconstructed top and antitop mass difference, $\Delta m^{\rm kin}_{t\bar{t}}$, an observable which is uniquely sensitive to CPT violation, to extract the first constraints on various coefficients responsible for it in the top-quark sector of the SME. Dedicated measurements of the kinematically reconstructed mass difference have been performed by DØ [13], CDF [14], ATLAS [15], and CMS [16, 17]. [^1] In Ref. [18], earlier DØ  and CDF top mass measurements [19, 20] were used to place a conservative upper limit on the top and antitop mass difference. All results are consistent with SM expectations within the experimental uncertainties, and may be translated into constraints on top quark coefficients for CPT violation.
@@
The kinematical mass difference for single event in terms of the SCF coefficients is given by
@@
To summarize, we have established the first model-independent sensitivity to CPT violation in the top-quark sector using LHC measurements of the top and antitop kinematical mass difference. The constraint on $b_{T}$ (13) is about two orders of magnitude stronger than those expected from single-top production [21]. We have also suggested dedicated analyses which would lead to constraints on $|b_{Z}|$ (23) and $|b_{X,Y}|$ (26) of the same order.
Regarding prospects, let us note that an analysis of the the entire Run-2 dataset (140 fb-1), assuming a factor of two improvement on the systematic uncertainty and taking into account the larger $\sqrt{s}=13\;{\rm TeV}$ center-of-mass energy, is expected to yield a sensitivity to $b_{T}$ at the 0.05 GeV level (with similar fractional improvements for $b_{X,Y,Z}$). Further accumulation of data will lead to a systematically dominated total uncertainty.
@@ Acknowledgments
AB and SM are supported in part through the NExT Institute and STFC Consolidated Grant No. ST/L000296/1. SM is also supported by the Knut and Alice Wallenberg foundation under the grant KAW 2017.0100. NS is supported in part by the Science and Technology Facilities Council (grant number ST/T006048/1). We thank V. A. Kostelecký for useful comments and discussions.

### v2

@@ Abstract
CPT invariance is the symmetry under the simultaneous transformations of charge conjugation (C), parity transformation (P), and time reversal (T). It is observed to be an exact symmetry of nature at the moment. According to the well-known CPT theorem [1, *Pauli:1955, *Luders:1957bpq], a local, unitary, and Lorentz invariant quantum field theory in Minkowski spacetime is CPT invariant. The CPT symmetry ensures that physical observables, including masses, lifetimes, magnetic moments and cross sections of any particle and its antiparticle are the same. It should also be stressed that the violation of Lorentz invariance does not necessarily imply the violation of CPT invariance.
Connecting theoretical predictions of CPT invariance violation to particle–antiparticle CPT tests has been carried out for all particle species of the Standard Model (SM) with exception of the top quark. This work addresses this gap, establishing the first sensitivity to top-sector CPT violation from top and antitop mass reconstructions in ATLAS and CMS Collaboration measurements.
Kostelecký and Potting demonstrated that spontaneous CPT violation in string theory may result in remnant observables (likely suppressed by the energy scale of this violation) which can be possibly tested in current experiments [4, *kp95]. The subsequent development of effective field theory (EFT) descriptions [6, *ck98, *cg97, *cg98], now referred to as the Standard Model Extension (SME), generated intense interest in testing CPT and Lorentz invariance across a variety of systems [10]. Within the SME framework, CPT violation necessarily implies Lorentz violation, while converse does not necessarily hold. Lorentz- and CPT-violating coefficients described by the SME can be viewed as generalized tensor background fields. In view of the discovery of the Higgs boson [11, 12] described by a homogeneous and isotropic scalar background with $SU(2)\times U(1)$ quantum numbers, searching for new physics in the form of additional backgrounds is well motivated.
In this study, our focus is on CPT-violating SME operators involving top-quark fields. We suggest using measurements of the kinematically reconstructed top and antitop mass difference, $\Delta m^{\rm kin}_{t\bar{t}}$, an observable which is uniquely sensitive to CPT violation, to extract the first constraints on various coefficients responsible for it in the top-quark sector of the SME. Dedicated measurements of the kinematically reconstructed mass difference have been performed by DØ [13], CDF [14], ATLAS [15], and CMS [16, 17]. [^1] In Ref. [18], earlier DØ  and CDF top mass measurements [19, 20] were used to place a conservative upper limit on the top and antitop mass difference. All results are consistent with SM expectations within the experimental uncertainties, and may be translated into constraints on top quark coefficients for CPT violation. We emphasize this approach is conceptually different from tests of Lorentz violation in the top sector [21, 22], which are insensitive to (leading-order) signatures of CPT violation [23].
@@
The kinematical mass difference for a single event in terms of the SCF coefficients is given by [^3] Note that in Eq. (15), we omit the helicity-dependent terms from Eq. (6) as they vanish under phase-space and time averages.
@@
To summarize, we have established the first model-independent sensitivity to CPT violation in the top-quark sector using LHC measurements of the top and antitop kinematical mass difference. The constraint on $b_{T}$ (13) is about two orders of magnitude stronger than those expected from single-top production [23]. We have also suggested dedicated analyses which would lead to constraints on $|b_{Z}|$ (23) and $|b_{X,Y}|$ (26) of the same order. Note that, although constraints on CPT-violating couplings involving $b$ and light quarks from $B$-meson and related oscillations are stringent, they do not directly constrain the third-generation $b$-quark coupling for CPT violation, which is proportional to $(a_{Q}+a_{D})_{33}$. For instance, in $B$-meson oscillations, the observable is governed by the difference $\Delta a=(a_{Q}+a_{D})_{33}-(a_{Q}+a_{D})_{22;\text{or};11}$ across generations [38], which could vanish if CPT violation is flavor universal. While existing results provide valuable indirect tests of CPT, our work offers a more direct approach focused on the $t$-quark coupling.
Regarding prospects, One might expect LHCb could offer complementary insights via $t\bar{t}$ asymmetry enhancements in the forward kinematic region. However, the quark-antiquark versus gluon-initiated top pair production is not relevant to the observable and analysis proposed in our work, and LHCb cannot fully reconstruct the top-antitop system for the mass difference determination. Therefore, we do not expect complementary insights from LHCb top-quark studies for this specific study. Let us note that an analysis of the the entire Run-2 dataset (140 fb-1), assuming a factor of two improvement on the systematic uncertainty and taking into account the larger $\sqrt{s}=13\;{\rm TeV}$ center-of-mass energy, is expected to yield a sensitivity to $b_{T}$ at the 0.05 GeV level (with similar fractional improvements for $b_{X,Y,Z}$). Further accumulation of data will lead to a systematically dominated total uncertainty.
@@ Acknowledgments
AB and SM are supported in part through the NExT Institute and STFC Consolidated Grant No. ST/L000296/1. SM is also supported by the Knut and Alice Wallenberg foundation under the grant KAW 2017.0100. AB acknowledges support from the Leverhulme Trust project MONDMag (RPG-2022-57). NS is supported in part by the Science and Technology Facilities Council (grant numbers ST/T006048/1 and ST/Y004418/1). EL acknowledges support from the Indiana University Center for Spacetime Symmetries and CERN. We thank V. A. Kostelecký for useful comments and discussions.